0.0.1 KAGRA Project

図 1: Conceptual picture of KAGRA. A 3-km ’L’- shaped interferometer is placed in the Kamioka un- derground site. KAGRA, previously called LCGT (Large Cryogenic Gravitational-Wave Antenna), is a 3-km laser interferometric gravitational wave antenna built at Kamioka underground site in Japan. With characteristic features of stable environment of the underground site and cryogenic technologies, KAGRA will be one of the most sensitive gravitational-wave telescopes in the world when it starts observation run in the year 2017. For a coalescence of binary neutron stars, KAGRA will have an observation range more than 200 Mpc. With this range, it is estimated that KAGRA will detect multiple gravitational-wave events in oneyear observation run. Along with the other telescopes, such as aLIGO in U.S.A. and Advanced VIRGO in Europe, KAGRA will open a new field of gravitational-wave astronomy.

KAGRA project was started in 2010 with a support by MEXT (Ministry of Education, Culture, Sports, Science, and Technology) of Japan. KAGRA is hosted by Institute of Cosmic-Ray Research (ICRR), the University of Tokyo, and is co-hosted by High Energy Accelerator Research Organization (KEK) and National Astronomical Observatory of Japan (NAOJ). KAGRA is built by a collaboration of more than 50 institutions and universities, or more than 150 research individuals. Among them, RESCEU has been playing an important role since the initial-design phase of KAGRA. Now, KAGRA is in the first construction phase called iKAGRA (initial KAGRA), in which a simplified 3-km interferometer will be constructed in the underground facility. The tunnel and facility will be ready in FY2013, and the first observation run will start with simplified configuration in 2015. After that, KAGRA will be upgraded to bKAGRA (baseline KAGRA). The interferometer configuration will be upgraded to RSE (Resonant-Sideband Extraction), and cryogenic system will be implemented. The first observation run with the full configuration is planned by the end of 2017.

図 2: Construction status of KAGRA. Left: The excavated hall for the Y-end room at the Kamioka underground site. Middle: A Facility to test the installation procedure for the 3-km arm ducts. About 70% of the arm duct has been completed. Right: Cryostats to house the KAGRA main mirrors.